Reduce the following quadratic form to standard form and find its principal axes:

4x^2-4xy+y^2


1
Expert's answer
2021-08-18T14:15:56-0400

"4x^2-4xy+y^2\\
\begin{pmatrix}
 4 & -2 \\
 -2 & 1
\end{pmatrix}\\
\mathrm{The\:eigenvalues\:of}\:A\:\mathrm{are\:the\:roots\:of\:the\:characteristic\:equation}\:\det \left(A-λ\:I\right)=0\\
λ=5,\:λ=0\\"

Therefore the orthogonal canonical reduction of the quadratic form is

"𝑄 = (𝑥′ 𝑦′) \begin{pmatrix}
 0 & 0 \\
 0 & 5
\end{pmatrix}\begin{pmatrix}
 x \\
 y
\end{pmatrix} = 5y'^2"

Principal axes at "λ=0\\"

The reduced row echelon form of the matrix is "\begin{bmatrix}
 1 & 2\\
 0 & 0
\end{bmatrix}"

To find the null space, solve the matrix equation"\begin{bmatrix}
 1 & 2\\
 0 & 0
\end{bmatrix} \begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}= \begin{bmatrix}
 0\\
 0 
\end{bmatrix}"

If we take "x_\n\n2\n\n=t\n\nx_2=t," then "x_1\u200b=\u22122tx_1\u200b=\u22122t" .

"Thus, x⃗ =\begin{bmatrix}
 -2t \\
 t
\end{bmatrix}= \begin{bmatrix}
 -2 \\
 1
\end{bmatrix} t"

"|x|= \sqrt{(-2)^2+1^2}=\sqrt5"

"\frac{1}{\sqrt5}\begin{pmatrix}
 -2 \\
 1
\end{pmatrix}"

Principal axes at "\u03bb=5"

The reduced row echelon form of the matrix is "\begin{bmatrix}
 1 & -0.5\\
 0 & 0
\end{bmatrix}"

To find the null space, solve the matrix equation"\begin{bmatrix}
 1 & -0.5\\
 0 & 0
\end{bmatrix} \begin{bmatrix}
 x_1 \\
 x_2
\end{bmatrix}= \begin{bmatrix}
 0\\
 0 
\end{bmatrix}"

If we take "x_\n\n2\n\n=t\n\nx_2=t," then "x_1\u200b=\u22122tx_1\u200b=\u22122t" .

"Thus, x⃗ =\begin{bmatrix}
 -0.5t \\
 t
\end{bmatrix}= \begin{bmatrix}
 -0.5 \\
 1
\end{bmatrix} t"

"|x|= \sqrt{(-0.5)^2+1^2}=\frac{\sqrt5}{2}"

"\frac{2}{\sqrt5}\begin{pmatrix}
 0.5 \\
 1
\end{pmatrix}"